# The Noisy Path from Source to Citation: Measuring How Scholars Engage with Past Research

### v2

@@ Abstract
Academic citations are widely used for evaluating research and tracing knowledge flows. Such uses typically rely on raw citation counts and neglect variability in citation types. In particular, citations can vary in their fidelity as original knowledge from cited studies may be paraphrased, summarized, or reinterpreted, possibly wrongly, leading to variation in how much information changes from cited to citing paper. In this study, we introduce a computational pipeline to quantify citation fidelity at scale. Using full texts of papers, the pipeline identifies citations in citing papers and the corresponding claims in cited papers, and applies supervised models to measure fidelity at the sentence level. Analyzing a large-scale multi-disciplinary dataset of approximately 13 million citation sentence pairs, we find that citation fidelity is higher when authors cite papers that are 1) more recent and intellectually close, 2) more accessible, and 3) the first author has a lower H-index and the author team is medium-sized. Using a quasi-experiment, we establish the "telephone effect" – when citing papers have low fidelity to the original claim, future papers that cite the citing paper and the original have lower fidelity to the original. Our work reveals systematic differences in citation fidelity, underscoring the limitations of analyses that rely on citation quantity alone and the potential for distortion of evidence.
@@ 1 Introduction
Citations are an essential part of scholarly communication. Given the assumption that authors cite prior work to acknowledge their intellectual debts to it Kaplan (1965); Harter (1992), citations are widely used to measure scientific contributions and trace the flows of knowledge. In practice, analysts and scholars tend to rely on raw citation counts Waltman (2016), implicitly assuming that citations are similar in kind, or at least that the variation is minor.
@@
Citations may paraphrase, summarize, or even misrepresent original knowledge, thereby leading to varying degrees of fidelity between the original and the citing claims. This variability in citation fidelity is not merely a theoretical concern; it has tangible consequences for the transmission of scientific knowledge. For instance, a short letter published in 1980 documenting opioid use in an inpatient sample Porter and Jick (1980) was subsequently cited uncritically as evidence supporting the safe use of opioid prescriptions Deyo et al. (2015). This letter had been cited over 600 times by 2017, with fewer than 20% of the citing publications acknowledging the critical information of the original context Leung et al. (2017). Such citations collectively contributed to a misleading narrative of low addiction risk, which played a role in the subsequent opioid epidemic West and Bergstrom (2021).
@@
Our analysis reveals several patterns in citation fidelity: fidelity is higher when authors cite papers that are more recent and intellectually close and more accessible; within teams, citation fidelity is highest for medium-sized teams, negatively correlated with the seniority of first authors, and shows no significant correlation with the seniority of last authors. Additionally, we identify a “telephone effect” of intermediary citations. Specifically, using a quasi-causal experimental setting, we find that the fidelity of citations is lower when the authors cite the original paper and an intermediary one that cites the original as well. Moreover, fidelity is positively correlated with the fidelity of the citation from the intermediary paper to the original paper citation. This telephone effect suggests that indirect engagement with original sources can result in information loss.
@@ 2.1 Citation classification
The normative theory of citation Kaplan (1965); Merton (1979) posits that authors cite papers to acknowledge their intellectual debts to them; thus, citations should indicate a direct semantic relationship or relevance between citing and cited works Garfield (1979). However, citation practices are not always accurate or rigorous, with issues such as misquotation, miscitation, and misrepresentation frequently observed Harzing (2002); Horbach et al. (2021); Cobb et al. (2024).
@@
A growing body of research has examined the relationship between citation and cited papers, revealing significant variations in citation fidelity West and Bergstrom (2021); Broadus (1983); Greenberg (2009). Studies have investigated cases of citations deviating from the original findings with low fidelity and have revealed that between 1 in 5 and 1 in 10 citations are used to support claims incongruous with the results of the cited paper Todd et al. (2010). For example, Cobb et al. (2024) examined 3,347 citing claims in top psychology journals and found that approximately 19% either omitted crucial nuances or entirely misrepresented the claims of previous research. Greenberg (2009) showed how hypotheses can be transformed into accepted facts through affirmative citation chains lacking direct supporting evidence in biomedical research. These findings reveal the variability in citation fidelity and low fidelity in citation can contribute to the noisiness of the knowledge transmission pipeline. However, these studies are often limited in scale and reliant on labor-intensive manual verification which requires substantial domain expertise, leaving gaps in our understanding of citation fidelity across broader academic discourse.
@@
Existing citation text studies have scaled citation classification by automating the identification of citation intent or function, primarily leveraging the immediate context surrounding the reference or the metadata of the citing paper. Cohan et al. (2019); Lauscher et al. (2022); Pham and Hoffmann (2003); Li et al. (2013); Pride and Knoth (2017); Valenzuela-Escarcega et al. (2015); Teufel et al. (2006); Baig et al. (2021). For example, Jurgens et al. (2018) classified the broad thematic functions of citations into six categories including background, motivation and uses. While these studies have expanded the scale of citation analysis, they still primarily focus on the citing paper, with limited efforts to use text of the cited paper Luu et al. (2021), particularly for assessing citation fidelity. Moreover, measuring citation fidelity extends beyond standard semantic textual similarity and requires assessing whether the citation faithfully conveys the same scientific information as the original text Wright et al. (2022). This task involves capturing more nuanced aspects of similarity, such as abstraction level, uncertainly, and contextual appropriateness August et al. (2020).
@@ 2.2 Author’s engagement with literature
A common criticism of using citations to indicate impact or trace knowledge flow is that authors may cite without directly engaging with the original source or being influenced by its content. There is widespread allegation that scientists may not fully read or only superficially review, the works they reference Simkin and Roychowdhury (2002); Jergas and Baethge (2015). Kaplan (1965) famously raised concerns about this practice, asking: “How often are the works of others cited without having been read carefully? How often are citations simply lifted from the bibliography in someone else’s work without either reading or giving credit to the man who did the original search of the literature?”
@@
However, less attention has been given to how engagement is reflected in subtle textual transformations, where original scientific information is reinterpreted or altered in nuanced ways with low fidelity. Studies suggest that such transformations with low fidelity may pose deeper challenges to knowledge transmission, contributing to issues such as the perpetuation of dominant discourses Mizruchi and Fein (1999) or even the propagation of misinformation in science West and Bergstrom (2021).
@@ 3.1 Dataset
In this study, we employ one of the most comprehensive academic datasets available: The Semantic Scholar Open Research Corpus (S2ORC) Lo et al. (2020). S2ORC is a general-purpose corpus containing metadata for 136 million scholarly articles spanning different fields. This dataset also provides access to the full text of articles, thus allowing us to analyze how authors make citations and identify which parts of the original text are being referenced on a large scale. After filtering out non-English articles and publications without full text, we obtained a dataset of approximately 42 million papers.
@@ 3.2 Reporting sentences from citing papers
Identifying single-source citations We first locate the in-text citation by identifying the sentence where the reference appears. Lauscher et al. (2022) on citation context show references to prior literature can sometimes span multiple sentences, though 83% of citation context occur within a single sentence. Given that most citations fall within this single-sentence window, we restrict our analysis to single-sentence citations for consistency and precision.
As citations occur at varying levels of aggregation and granularity Cronin (1994), especially when multiple sources are cited together, we focus on citations that explicitly report claims from a single source. To further refine this selection, we apply a rule-based filter, considering only parenthetical citations enclosed in [] or () at the sentence’s end. This structured approach helps isolate the core reporting function of citations and minimizes confounding factors, enhancing the precision of our analysis. See Appendix Table 2 for valid case examples.
Identifying background citations Citations can serve multiple functions beyond reporting claims from the original paper. Previous research has developed classification schemes for citations based on their roles within an argument. Among these, background citations are widely recognized as a category that provides essential context and relevant information for the domain under discussion Teufel et al. (2006); Cohan et al. (2019); Lauscher et al. (2022). These citations often set the stage for a study by informing readers about existing knowledge, highlighting research gaps, and establishing the context for the study’s purpose Jurgens et al. (2018).
Leveraging their framework, we focus on background citations as a means to further identify reporting sentences. We constructed a more comprehensive dataset by integrating annotated data from two prominent sources for citation function classification: SciCite Cohan et al. (2019) and MultiCite Lauscher et al. (2022), both of which annotate background as a citation use. The combined dataset includes 27,052 instances, with 11,635 (43%) labeled as background. We fine-tuned a SciBERT model Beltagy et al. (2019) based on the integrated dataset for a binary classification task to identify whether a sentence serves as a background citation. The model achieved an F1 score of 0.81 using an 80-20 train-test split, which shows its effectiveness in accurately detecting background citations.
@@ 3.3 Sentences of claims from cited papers
To match reporting citations with the corresponding claims in cited papers, we apply a sentence classifier to identify result or conclusion sentences within the full text of the cited publication. This classifier is fine-tuned from a RoBERTa model Liu et al. (2019) and is trained to categorize sentences into five categories: methods, background, objective, results, and conclusions. The training dataset consists of 200,000 paper abstracts from PubMed, which were self-labeled with these categories Canese and Weis (2013). The classifier achieved an F1 score of 0.92 on a held-out 10% sample (details in Wright et al., 2022). Using this approach, we filter approximately 30% of sentences of results or conclusions from the full text as findings to be matched with citing sentences.
4 Measuring Citation Fidelity
4.1 Measuring sentence-level scientific information change
For this purpose, we employ a sentence-level scientific information change measure developed by Wright et al. (2022). This measure is specifically designed for evaluating scientific texts and was trained on an annotated dataset of scientific claims. It quantifies the extent to which two semantically similar sentences describe the same scientific claims and capture topic alignment as well as both nuance and accuracy in how claims are reported. The metric assigns a similarity score from 1 to 5, where: 1 indicates that the two sentences describe completely different claims, and 5 indicates that the sentences express identical claims (see Appendix Table 3).
Upper-bound approach Assessing fidelity between citation pairs is challenging because authors typically cite entire publications rather than specific sentences, making it difficult to determine which exact claim the citing sentence refers to. This indirect citation style complicates the task of pairing citing and cited claims.
@@ 4.1 Measuring sentence-level scientific information change
Table 1: Examples of citing and matched result sentences with fidelity scores. The first pair from Applewhite et al. (2022) and Ruan et al. (2018) shows high fidelity as both sentences discuss prenatal exposure to music and ASD, despite wording differences. The second pair from Canario et al. (2021) and Specht (2020) shows lower fidelity due to differing focuses on rs-fMRI limitations.
@@ 4.2 Interpreting citation fidelity
By extracting reporting citations and corresponding claims in cited sources, we generate 13 million citation pairs. A random sample of fidelity score shows that the majority cluster around 3.5, following a roughly normal distribution ranging from 1.0 to 5.0 (See Appendix Figure 6). This suggests that most citations exhibit moderate fidelity, with fewer cases of extremely high or low similarity. While the majority of citations maintain a reasonable level of fidelity, the presence of extreme values underscores the variability in how accurately authors represent cited claims. Moderate fidelity scores could reflect a balance between accurate representation and paraphrasing to fit the citing paper’s context, whereas lower scores might signal potential misrepresentation or information loss.
@@ 5 Higher Fidelity is Associated with Closer Proximity
Experimental setup We employ a standard regression analysis framework, with the model specified below. Citation fidelity is the dependent variable (F). This approach examines key factors influencing fidelity, including accessibility, temporal and disciplinary proximity, and team dynamics, while controlling for other relevant attributes (See Appendix Table 4 for more details on all related variables).  [^3] Key variables are defined as follows: Team dynamics is proxied using Team.Size and Author.Seniority, where Team.Size represents the number of authors in the team, and Author.Seniority represents the experience of the authors using their H-index. Accessibility is proxied by Open.Access, a binary variable indicating whether the cited paper is publicly accessible. Temporal and disciplinary proximity are proxied using Publication.Interval, Self.Citation, and With.Field, where Publication.Interval represents the number of years between the cited and citing papers, Self.Citation indicates whether the two papers share at least one author, and With.Field indicates whether they belong to the same discipline.
@@
Results Figures 2, 3, and 4 present coefficient estimates from regression analyses, controlling for potential confounders. Error bars represent standard errors (See Figure 7 for more regression results).

### v3

@@ Abstract
Academic citations are widely used for evaluating research and tracing knowledge flows. Such uses typically rely on raw citation counts and neglect variability in citation types. In particular, citations can vary in their fidelity as original knowledge from cited studies may be paraphrased, summarized, or reinterpreted, possibly wrongly, leading to variation in how much information changes from cited to citing paper. In this study, we introduce a computational pipeline to quantify citation fidelity at scale. Using full texts of papers, the pipeline identifies citations in citing papers and the corresponding claims in cited papers, and applies supervised models to measure fidelity at the sentence level. Analyzing a large-scale multi-disciplinary dataset of approximately 13 million citation sentence pairs, we find that citation fidelity is higher when authors cite papers that are 1) more recent and intellectually close, 2) more accessible, and 3) the first author has a lower H-index and the author team is medium-sized. Using a quasi-experiment, we establish the “telephone effect”—when citing papers have low fidelity to the original claim, future papers that cite the citing paper and the original have lower fidelity to the original. Our work reveals systematic differences in citation fidelity, underscoring the limitations of analyses that rely on citation quantity alone and the potential for distortion of evidence.
@@ 1 Introduction
Citations are an essential part of scholarly communication. Given the assumption that authors cite prior work to acknowledge their intellectual debts to it Kaplan (1965); Harter (1992), citations are widely used to measure scientific contributions and trace the flows of knowledge. In practice, analysts and scholars tend to rely on raw citation counts Waltman (2016), implicitly assuming that citations are similar in kind, or that any variation is minor.
@@
Citations may paraphrase, summarize, or even misrepresent original knowledge, thereby leading to varying degrees of fidelity between the original and the citing claims. This variability in citation fidelity is not merely a theoretical concern; it has tangible consequences for the transmission of scientific knowledge. For instance, a short letter published in 1980 documenting opioid use in an inpatient sample Porter and Jick (1980) was subsequently cited uncritically as evidence supporting the safe use of opioid prescriptions Deyo et al. (2015). This letter had been cited over 600 times by 2017, with fewer than 20% of citations acknowledging the critical information of the original context Leung et al. (2017). Such citations collectively contributed to a misleading narrative of low addiction risk, which played a role in the subsequent opioid epidemic West and Bergstrom (2021).
@@
Our analysis reveals several patterns in citation fidelity: fidelity is higher when authors cite papers that are more recent and intellectually close and more accessible; within teams, citation fidelity is highest for medium-sized teams, negatively correlated with the seniority of first authors, and shows no significant correlation with the seniority of last authors. Additionally, we identify a “telephone effect” of intermediary citations. Specifically, using a quasi-causal experimental setting, we find that the fidelity of citations is lower when the authors cite the original paper and an intermediary one that cites the original as well. Moreover, citation fidelity is positively correlated with the fidelity of the intermediary paper’s citation to the original paper. This telephone effect suggests that intermediary engagement with original sources can result in information loss. [^3] The codes and models for this paper are available at https://github.com/hongcchen/citation.
@@ 2 Motivation and Related Works
Citations are a widely used bibliometric tool for evaluating the performance of various actors in the scientific community. To better understand citation behavior, researchers have increasingly examined the content of citations, especially the linguistic and contextual features of citing text, not only to infer authors’ intent and citations’ function, but also to interrogate a core assumption: to what extent do citations reflect genuine scientific influence on the citing authors and their work?
@@ 2.1 Citation classification
The normative theory of citation Kaplan (1965); Merton (1979) posits that authors cite papers to acknowledge their intellectual debts to them; thus citations should indicate a direct semantic relationship or relevance between citing and cited works Garfield (1979). However, citation practices are not always accurate or rigorous, with issues such as misquotation, miscitation, and misrepresentation frequently observed Harzing (2002); Horbach et al. (2021); Cobb et al. (2024).
@@
A growing body of research has examined the relationship between citation and cited papers, revealing significant variations in citation fidelity West and Bergstrom (2021); Broadus (1983); Greenberg (2009); Zhang and Abernethy (2024). Studies have investigated cases of citations deviating from the original findings with low fidelity and have revealed that between 1 in 5 and 1 in 10 citations are used to support claims incongruous with the results of the cited paper Todd et al. (2010). For example, Cobb et al. (2024) examined 3,347 citing claims in top psychology journals and found that approximately 19% either omitted crucial nuances or entirely misrepresented the claims of previous research. Greenberg (2009) showed how hypotheses can be transformed into accepted facts through affirmative citation chains lacking direct supporting evidence in biomedical research. These findings reveal the variability in citation fidelity and low fidelity in citation can contribute to the noisiness of the knowledge transmission pipeline. However, these studies are often limited in scale and reliant on labor-intensive manual verification which requires substantial domain expertise, leaving gaps in our understanding of citation fidelity across broader academic discourse.
@@
Existing citation text studies have scaled citation classification by automating the identification of citation intent or function, primarily leveraging the immediate context surrounding the reference or the metadata of the citing paper Cohan et al. (2019); Lauscher et al. (2022); Pham and Hoffmann (2003); Li et al. (2013); Pride and Knoth (2017); Valenzuela-Escarcega et al. (2015); Teufel et al. (2006); Baig et al. (2021). For example, Jurgens et al. (2018) classified the broad thematic functions of citations into six categories including background, motivation, uses, extension, comparision or contrast and future. While these studies have expanded the scale of citation analysis, they still primarily focus on the citing paper, with limited efforts to use text of the cited paper Luu et al. (2021), particularly for assessing citation fidelity.
Moreover, measuring citation fidelity extends beyond standard semantic textual similarity and requires assessing whether the citation faithfully conveys the same scientific information as the original text Wright et al. (2022). While there are potential alternative metrics from related areas such as fact-checking and semantic similarity Thorne et al. (2018); Ganitkevitch et al. (2013), they are typically designed for different goals (e.g., truth verification, general paraphrase accuracy) and do not directly address fidelity in the context of scientific claims. This task of measuring fidelity requires capturing more nuanced aspects of similarity, such as abstraction level, uncertainty, and contextual appropriateness August et al. (2020).
@@ 2.2 Author’s engagement with literature
A common criticism of using citations to indicate impact or trace knowledge flow is that authors may cite without directly engaging with the original source or being influenced by its content. There is widespread allegation that scientists may not fully read or only superficially review the works they reference Simkin and Roychowdhury (2002); Jergas and Baethge (2015). Kaplan (1965) famously raised concerns about this practice, asking: “How often are the works of others cited without having been read carefully? How often are citations simply lifted from the bibliography in someone else’s work without either reading or giving credit to the man who did the original search of the literature?”
@@
However, less attention has been paid to how authors’ engagement is reflected in subtle textual signals—specifically, how they report original findings in their citations. An author’s level of engagement may be revealed in how they reinterpret or modify original scientific information, whereas limited understanding or familiarity can result in lower alignment or fidelity. Studies suggest that such low-fidelity citations may contribute to broader issues in knowledge transmission, such as the perpetuation of dominant discourses Mizruchi and Fein (1999) or even the propagation of misinformation in science West and Bergstrom (2021).
@@ 3.1 Dataset
In this study, we employ one of the most comprehensive academic datasets available: The Semantic Scholar Open Research Corpus (S2ORC; Lo et al., 2020). S2ORC is a general-purpose corpus containing metadata for 136 million scholarly articles spanning different fields. This dataset also provides access to the full text of articles, thus allowing us to analyze how authors make citations and identify which parts of the original text are being referenced on a large scale. After filtering out non-English articles and publications without full text, we obtained a dataset of approximately 42 million papers.
@@ 3.2 Reporting sentences from citing papers
Identifying single-source citations
We first locate the in-text citation by identifying the sentence where the reference appears, setting the citation context window to be a single sentence. Li et al. (2022) find that many citations tend to correspond to a very localized piece of the cited document typically ranging from roughly a short clause to spans of one or two sentences. Lauscher et al. (2022) also report that while references to prior literature can sometimes span multiple sentences, 83% of citation contexts occur within a single sentence. Given that most citations fall within this single-sentence window, we restrict our analysis to single-sentence citations for consistency and precision.
As citations occur at varying levels of aggregation and granularity Cronin (1994), especially when multiple sources are cited together, we focus on citations that explicitly report claims from a single source. To further refine this selection, we apply a rule-based filter, considering only parenthetical citations enclosed in [] or () at the sentence’s end. This structured approach helps isolate the core reporting function of citations and minimizes confounding factors, enhancing the precision of our analysis. See Appendix Table 2 for valid examples of citing sentences.
Identifying background citations
Citations can serve multiple functions beyond reporting claims from the original paper. Previous research has developed classification schemes for citations based on their roles within an argument. Among these, background citations are widely recognized as a category that provides essential context and relevant information for the domain under discussion Teufel et al. (2006); Cohan et al. (2019); Lauscher et al. (2022). These citations often set the stage for a study by informing readers about existing knowledge, highlighting gaps, and establishing the context for the study’s purpose Jurgens et al. (2018).
Leveraging their framework, we focus on background citations as a means to further identify reporting sentences. We construct a more comprehensive dataset by integrating annotated data from two prominent sources for citation function classification: SciCite Cohan et al. (2019) and MultiCite Lauscher et al. (2022), both of which annotate background as a citation function. The combined dataset includes 27,052 instances, with 11,635 (43%) labeled as background. We fine-tune a SciBERT model Beltagy et al. (2019) based on the integrated dataset for a binary classification task to identify whether a sentence serves as a background citation. The model achieves an F1 score of 0.81 using an 80-20 train-test split, which shows its effectiveness in accurately detecting background citations.
@@ 3.3 Sentences of claims from cited papers
To match reporting citations with the corresponding claims in cited papers, we follow the definition that background citations typically refer to relevant findings or essential context from the cited work. Therefore, we focus on sentences of claims, which are operationalized as those labeled as “results” or “conclusions” based on the discourse categories defined by Dernoncourt and Lee (2017). We apply a sentence classifier to identify result or conclusion sentences within the full text of the cited publication. This classifier is fine-tuned from a RoBERTa model Liu et al. (2019) and is trained to categorize sentences into five categories: methods, background, objective, results, and conclusions. The training dataset consists of 200,000 paper abstracts from PubMed, which were self-labeled with these categories Canese and Weis (2013). The classifier achieved an F1 score of 0.92 on a held-out 10% sample (details in Wright et al., 2022). Using this approach, we filter approximately 30% of sentences of results or conclusions from the full text as claims to be matched with citing sentences.
@@
Table 1: Examples of citing and matched sentences with fidelity scores. The first pair from Applewhite et al. (2022) and Ruan et al. (2018) shows high fidelity as both sentences discuss prenatal exposure to music and ASD, despite wording differences. The second pair from Canario et al. (2021) and Specht (2020) shows lower fidelity due to differing focuses on rs-fMRI limitations. The second pair also reflects an interesting bibliographic inconsistency: the source is cited as 2019 though the paper was published in 2020.
4 Measuring Citation Fidelity
Assessing citation fidelity requires measuring how faithfully a citation reflects the original findings of the cited work. This involves first identifying the corresponding claims in the cited text, and then analyzing how the citation relates to that claim. Both are challenging due to the nuanced nature of scientific information and the absence of explicit sentence-level referential links in citations. To address this, we employ a sentence-level measure specifically designed for scientific findings, along with necessary approximations.
4.1 Measuring sentence-level scientific information change
For this purpose, we employ a sentence-level scientific information change measure developed by Wright et al. (2022). This measure quantifies the extent to which two semantically similar sentences describe the same scientific claims and capture topic alignment as well as both nuance and accuracy in how claims are reported. It is specifically designed for evaluating scientific texts and is well-aligned with our task of citation fidelity, with minimal domain or task shift.
The metric assigns a similarity score from 1 to 5, where 1 indicates that the two sentences describe completely different claims, and 5 indicates that the sentences express identical claims (see Appendix Table 3). The model is fine-tuned from MPNet Song et al. (2020) on a human-annotated dataset (crowdsourced and expert-reviewed) and trained to minimize the cross-entropy over sentence embedding dot products. It achieves a mean square error of 0.489 and a Pearson correlation of 76.48 in predicting the scientific information change (details in Wright et al., 2022).
Upper-bound approach
Pairing the reporting citation and the corresponding claim is challenging because authors typically cite the entire publication rather than referring to specific sentences, making it difficult to determine which exact claim the citing sentence refers to. The absence of explicit sentence-level referential links complicates the task of assessing citation fidelity.
@@ 4.2 Interpreting citation fidelity
By extracting reporting citations and corresponding claims in cited sources, we generate 13 million citation pairs. A random sample of fidelity score shows that the majority cluster around 3.5, following a roughly normal distribution ranging from 1.0 to 5.0 (See Appendix Figure 6). This suggests that most citations exhibit a reasonable level of fidelity, though the presence of fewer extreme high and low values shows the variability in how accurately authors represent cited claims. Moderate fidelity scores could reflect a balance between accurate representation and paraphrasing to fit the citing paper’s context, whereas lower scores might signal information loss or potential misrepresentation.
@@ 5 Higher Fidelity is Associated with Closer Proximity
Experimental setup
We employ a standard regression analysis framework, with the model specified below. Citation fidelity is the dependent variable (F). Key factors such as accessibility, proximity, and team dynamics are included along with other relevant attributes (See Appendix Table 4 for more details on all variables).
@@
Results
Figures 2, 3, and 4 present coefficient estimates from regression analyses, controlling for potential confounders. The regression coefficients are mapped back to the original 0–5 fidelity scale by adding back the intercept, in order to visualize the results in a more interpretable way. Error bars represent standard errors (See Figure 7 for more regression results).  [^4] Key variables are defined as follows: Team dynamics is proxied using Team.Size and Author.Seniority, where Team.Size represents the number of authors in the team, and Author.Seniority represents the experience of the authors using their H-index. Accessibility is proxied by Open.Access, a binary variable indicating whether the cited paper is publicly accessible. Temporal and disciplinary proximity are proxied using Publication.Interval, Self.Citation, and With.Field, where Publication.Interval represents the number of years between the cited and citing papers, Self.Citation indicates whether the two papers share at least one author, and With.Field indicates whether they belong to the same discipline.
First, our findings reveal that proximity—whether in terms of accessibility, disciplinary context, temporal closeness, or author familiarity—plays a critical role in shaping citation fidelity. One strong piece of evidence is that self-citations, which is arguably a straightforward indicator of engagement as authors are naturally more familiar with their own work, exhibit higher fidelity scores than non-self-citations. This supports the validity of our fidelity measure, as it reflects the expected pattern: authors who engage more deeply with cited work tend to cite with higher fidelity.
Similarly, open-access publications exhibit higher fidelity compared to non-open-access publications, likely because unrestricted access to the original sources allows more direct engagement and accurate representation of cited work Nagaraj et al. (2020). Without access to the full content of the cited work, authors may only partially examine the original material, relying on overviews or others’ interpretations. Citation fidelity is higher for within-field citations compared to cross-field citations. This is likely because authors are more familiar with the methodologies, terminologies, or conceptual frameworks of their own disciplines, and are able to represent without much information change. In contrast, cross-field citations, while essential for interdisciplinary knowledge transfer Thilakaratne et al. (2018), may introduce additional information loss due to differences in their research norms, jargon, or methodologies.